SEPARATION AND CONVERSION OF NUCLEAR SPIN ISOMERS OF ETHYLENE

\underline{Z.-D. SUN}, Department of Physics, Toyama University, 930-8555, Japan;Department of Physics, Yantai University, 264005, P. R. China; K. TAKAGI, S. YAMADA, H. TANAKA, and F. MAT-SUSHIMA, Department of Physics, Toyama University, 930-8555, Japan.

The separation and conversion of nuclear spin isomers with A_{g}, B_{1g}, B_{2u}, and B_{3u} symmetry of normal ethylene (C_{2}H_{4}) have been investigated by a newly assembled experimental setup in the first case studied so far for a planar, asymmetrictop polyatomic molecule that has four nuclear spin isomers with the D_{2h} symmetry group. One of the nuclear spin isomers of C_{2}H_{4} is first separated using the 10P 44 CO_{2} laser line in the method of Light-Induced Drift^{a}. The equilibration decay curves of the spin isomers are then measured using a second probe CO_{2} laser to monitor absorption spectral line intensities as a function of time. The conversion rates between isomers have been determined by fitting these decay curves to an exponential function. The experimental results and their interpretations will be presented.

aF.Kh. Gelmukhanov, and A.M. shalagin, JETP Lett. 29, 711 (1979).